b1az34:

These are the first sentences in the short story, "The Gift of the Magi," by O. Henry: "One dollar and eighty-seven cents. That was all. And sixty cents of it was in pennies." What are possible numbers of coins to make $1.87? Use the equation that represents the value of the coins to complete the table.

b1az34:

b1az34:

what the formula or how do i solve this

daddywolfhotstuff:

Each row is read from left to right not down, that's what messed me up. All it wants you to do is fill in the blanks to make each row equal 187. It's a tricky table. q+d+n+p = 187. q=25, d=10, n=5, p=1. So for the first row, it's 4q + 6d + 5n + 2p = 187.

daddywolfhotstuff:

is that good

b1az34:

i got confused on how 4+6+5+2= 187

daddywolfhotstuff:

ohhhhh understanable

b1az34:

do i subsisted q=25, d=10, n=5, p=1

b1az34:

for each row or how?

daddywolfhotstuff:

i think

b1az34:

ho ok thanks for clearing some of it up

daddywolfhotstuff:

hey we both on lvl 25

b1az34:

do i subsisted q=25, d=10, n=5, p=1 for each row

daddywolfhotstuff:

welcome anytime i like math

b1az34:

alr